OpenStudy (anonymous):

given f(x)= 6-2x and g(x) = 5x-12, f(g(x))= ? (show work)

OpenStudy (bibby):

plug g(x) into f(x) \(f((g(x))=6-2(f(x))=6-2(5x-12)\)